Robert has money in two savings accounts. One rate is 6% and the other is 14%. If he
has $600 more in the 14% account and the total interest is $186, how much is invested
in each savings account?

Answer by fractalier(6550)

Call the amount of money in the 6% account, x.
Then the amount in the 14% account is 600+x.
The set up is then
.06x + .14(600+x) = 186
.06x + 84 + .14x = 186
.20x = 102
x = $510 at 6%
600+x = $1110 at 14%